OpenStudy (anonymous):

Lincoln and Gemma are looking at the equation sqrt(2x - 3) = sqrt(x). Lincoln says that the solution is extraneous. Gemma says the solution is non-extraneous. Is Lincoln correct? Is Gemma correct? Are they both correct? Justify your response by solving this equation, explaining each step with complete sentences.

OpenStudy (anonymous):

@peachpi Please help me, haha. I'm stuck on this and have been for about 45 minutes.

OpenStudy (anonymous):

I think there should be a non-extraneous solution. I'd start by dropping the square roots to make the equation linear

OpenStudy (anonymous):

2x -3 = x x=3

OpenStudy (anonymous):

This is what I have so far: sqrt(2x - 3) = sqrt(x) 2x - 3 = x I used mathway to check what the answer should be and it's 3, so I checked that and it is sqrt(2(3) - 3) = 3 2(3) - 3 = 3 6 - 3 = 3 3 = 3

OpenStudy (anonymous):

What happens to the 2?

OpenStudy (anonymous):

It was multiplied by 3 in the simplification

OpenStudy (anonymous):

In the part that you did? 2x -3 = x x=3

OpenStudy (anonymous):

I was looking at your work. Where 2(3) - 3 turns into 6 - 3 on the left side

OpenStudy (anonymous):

In mine I subtracted x and added 3 to both sides.

OpenStudy (anonymous):

Okay, but it yours where does the 2 go? P.S. I'm sorry, I suck at math

OpenStudy (anonymous):

I subtracted the x on the right from 2x. 2x - x = x